# KnowGPT: Knowledge Injection for Large Language Models

### v4

@@ Abstract
Generative Large Language Models (LLMs), such as ChatGPT and GPT-4, offer interactive APIs that can answer common questions at the human-expert level. However, these models often give inaccurate responses when faced with questions requiring domain-specific or professional-specific knowledge not covered in their training corpus. Furthermore, many state-of-the-art LLMs are not open-source, making it challenging to inject knowledge with model APIs only. In this work, we introduce KnowGPT, a black-box knowledge injection framework for LLMs in question answering. KnowGPT leverages deep reinforcement learning (RL) to extract relevant knowledge from Knowledge Graphs (KGs) and use Multi-Armed Bandit (MAB) to construct the most suitable prompt for each question. Our extensive experiments on three benchmark datasets showcase that KnowGPT significantly enhances the existing methods. Notably, KnowGPT achieves an average improvement of 23.7% over ChatGPT and an average improvement of 2.9% over GPT-4. Additionally, KnowGPT attains a 91.6% accuracy on the OpenbookQA official leaderboard, which is comparable to human-level performance. The code will be open-sourced.
KnowGPT: Knowledge Injection for Large Language Models
Qinggang Zhang1*, Junnan Dong1*, Hao Chen1, Daochen Zha2, Zailiang Yu3, Xiao Huang1 1The Hong Kong Polytechnic University, 2Rice University, 3Zhejiang Lab {qinggangg.zhang,hanson.dong}@connect.polyu.hk, sundaychenhao@gmail.com daochen.zha@rice.edu, yuzl@zhejianglab.com, xiaohuang@comp.polyu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have surprised the world with their superior performance Kung et al. (2023); Zha et al. (2023), especially with the emergence of ChatGPT and GPT-4 OpenAI (2023). Nonetheless, LLMs are often criticized for their limited factual knowledge and propensity to produce hallucinations, wherein the model fabricates incorrect statements on tasks beyond their knowledge and perception Amaro et al. (2023); Shen et al. (2023); Gravel et al. (2023). Consider an ecological domain-specific question from OpenbookQA Mihaylov et al. (2018) in Figure 1. ChatGPT erroneously responds “energy” when asked about the portion of nutrients. This inaccuracy could stem from its potential lack of knowledge of carbs and their relationship to nutrients.
@@
Figure 1: A real-world question in OpenbookQA and a subgraph of ConceptNet. ChatGPT could effectively correct the answer given the scientific reasoning background in KG (blue: question concepts, red: candidate answers).
A promising avenue for addressing the above issue entails the integration of Knowledge Graphs (KGs) into LLMs. KGs Li et al. (2022a, b, 2023); Zhang et al. (2022a, 2023), such as Yago Suchanek et al. (2007), Freebase Bollacker et al. (2008), and ConceptNet Speer et al. (2017) represent relationships among real-world entities in a structured form as triples (head, relation, tail). The enormous factual knowledge stored in KGs holds the potential to significantly enhance the accuracy of LLMs’ responses. For instance, in Figure 1, ChatGPT could correct itself by leveraging the related background knowledge in ConceptNet Speer et al. (2017).
Recently, various white-box knowledge injection methods have been proposed to incorporate KGs into LLMs. These methods assume that we know everything about LLMs, such as model architectures and parameters. Many algorithms have been proposed to integrate KGs into LLMs Pan et al. (2023). Early studies directly concatenate the entities from KGs with textual sentences as the input to train LLMs based on cross-modal representation alignment Sun et al. (2020b); Liu et al. (2020). Later research incorporates KGs at an implicit level by either combining text representation and KG embedding with the attention mechanism Feng et al. (2020); Yasunaga et al. (2021); Dong et al. (2023a) or designing various fusion methods to integrate KGs and texts through a tailored encoder Zhang et al. (2019); Sun et al. (2020a).
Unfortunately, many state-of-the-art LLMs are confined to be close-source in practical applications. For instance, ChatGPT and GPT-4 exclusively grant access through their APIs, which means we can only retrieve model responses by submitting textual inputs, with model specifics inaccessible. This lack of access prevents us from employing the aforementioned white-box knowledge injection techniques. Even though white-box approaches could be applied to open-source LLMs, such as BLOOM Scao et al. (2022) and LLaMA Touvron et al. (2023), they will often incur significant computation costs due to updating model weights Liu et al. (2022). Thus, we ask: Can we develop a knowledge injection framework that can efficiently and effectively integrate KGs into LLMs with APIs only?
Achieving this goal is nontrivial because of two challenges in constructing model inputs (prompts). Identifying the most relevant knowledge is difficult. Real-world KGs often consist of millions of triples, whereas LLMs are typically restricted by limited input lengths (e.g., 2048 tokens for ChatGPT and 4096 tokens for GPT-4). Hence, careful selection of the most informative knowledge from KGs becomes essential. Effectively encoding KG knowledge is hard. It is observed that even minor variations in prompts conveying the same semantic meaning can yield drastically different responses from LLMs OpenAI (2023). As a result, a customized approach to encoding factual knowledge from extracted KGs for each question is often required to achieve the best performance.
@@
In this work, we propose KnowGPT, a knowledge injection framework for LLMs in question answering. To address challenge , we leverage deep reinforcement learning (RL) to extract paths from source entities mentioned in the question to the target entities within the potential answers. To encourage the agent to discover more informative paths, we devise a tailored reward scheme that promotes the reachability, context-relatedness, and conciseness of the extracted paths. Then, a policy network is trained to maximize the reward using training questions and applied to unseen questions. To tackle challenge , we introduce a prompt construction strategy based on Multi-Armed Bandit (MAB). Given several path extraction strategies and prompt templates, a MAB is learned to select the most effective combination for each question by balancing exploration and exploitation. The learned MAB is then applied to new questions to select path extraction strategies and prompt templates automatically. Our main contributions are:
Figure 2: The overall architecture of KnowGPT. Given the question context and answer choices, we retrieve a question-specific subgraph from the real-world KG. Knowledge Extraction is first dedicated to searching for the most informative and concise reasoning background subject to the context. Then the prompt translation module is optimized to prioritize the combination of knowledge and formats subject to the given question context.
  • Formally define the problem of balck-box knowledge injection for LLMs, which integrates KGs into LLMs with model APIs only.
  • Propose KnowGPT, a general knowledge injection model to capture and translate knowledge from Knowledge Graphs (KGs) into prompts.
  • Instantiate KnowGPT upon two real-world KGs, ConceptNet Speer et al. (2017) and USMLE Yasunaga et al. (2021), with ChatGPT APIs. KnowGPT outperforms the state-of-the-art baselines on three QA benchmarks by a large margin. Notably, KnowGPT attains a 91.6% accuracy on the OpenbookQA leaderboard, which is comparable to human performance.
2 Related Work
2.1 Integration of KGs and LLMs
Leveraging KG to assist LLMs for NLP tasks has been intensively studied. Colake Sun et al. (2020b) presents a unified graph that combines a word graph with the given context and KG. QA-GNN Yasunaga et al. (2021) takes this further and embeds the question context as an entity in the joint graph. Recent studies utilize the attention mechanism to incorporate KGs into LMs, thereby enhancing comprehension and reasoning processes. MHGRN Feng et al. (2020) iteratively refines the question embeddings from pre-trained LLMs via graph neural networks during reasoning. However, they are limited to adapt to LLMs since most existing LLMs are either closed-sourced or over-expensive to be effectively tuned.
To alleviate this issue, recent research has shown various ways to prompt the KG to LLMs. One heuristic way to incorporate KGs and LLMs is to inject triples as the input Pan et al. (2023). ERNIE3.0 Sun et al. (2021) takes triples as token sequences and straightforwardly appends them to the given sentences. Another group of work dedicate to designing various ways to retrieve the knowledge. Mindmap Wen et al. (2023) transforms the relevant knowledge into mindmaps and prompt the LLMs. ChatRule Luo et al. (2023) extracts multiple relational paths from KGs and prompts the LLMs with natural language sentences. CoK Wang et al. (2023) activates the LLMs’ comprehension through chain-of-knowledge which prompts triples step-by-step. However, they lack the prompt format design which is laborious to be expanded to more LLMs and task scenarios. In this work, we develop a knowledge injection framework that values both knowledge retrieval and prompt construction. It automatically prioritize relevant knowledge and prompts the LLMs with the the most suitable prompt formats.
@@ 3 Problem Statement
We formally define the problem of black-box knowledge injection for LLMs in complex question answering. We represent each question as a question context $\mathcal{Q}=\{\mathcal{Q}_{s},\mathcal{Q}_{t}\}$, where $\mathcal{Q}_{s}=\{e_{1},e_{2},...,e_{m}\}$ is a set of $m$ source entities, and $\mathcal{Q}_{t}=\{e_{1},e_{2},...,e_{n}\}$ is a set of $n$ target entities. Following prior work Feng et al. (2020); Yasunaga et al. (2022), $\mathcal{Q}_{s}$ is extracted by concept recognition, and we assume it is given in our problem. Similarly, each target entity in $\mathcal{Q}_{t}$ is extracted from a corresponding candidate answer. We denote an LLM as $f$, a real-world KG as $\mathcal{G}$, which consists of triples (head entity, relation, tail entity), denoted as $(h,r,t)$. In our setting, we only have access to the APIs of $f$. However, we can employ open-source lightweight language models (not $f$), like Bert-Base Kenton and Toutanova (2019), to obtain text embeddings. Using the above notations, we describe our problem below. Given a question context $\mathcal{Q}$, an LLM $f$, and a KG $\mathcal{G}$, we aim to learn a prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$, which generates a prompt x that incorporates the context of $\mathcal{Q}$ and the factual knowledge in $\mathcal{G}$, such that the prediction of the LLM $f(\textbf{x})$ can output the correct answers for $\mathcal{Q}$.
@@ 4 KnowGPT Framework
Learning the prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$ involves two challenges, i.e., what knowledge should be used in $\mathcal{G}$, and how to translate the structured knowledge into prompts. To address these challenges, we present KnowGPT, which extracts subgraphs (paths) with deep RL and then constructs the prompt with MAB. An overview of our framework is shown in Figure 2.
@@ 4.1 Knowledge Extraction with Deep Reinforcement Learning
Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{\text{sub}}$ that contains all the source entities $\mathcal{Q}_{s}$, target entities $\mathcal{Q}_{t}$, and their neighbors. An ideal subgraph $\mathcal{G}_{\text{sub}}$ is expected to have the following properties: $(i)$ $\mathcal{G}_{\text{sub}}$ encompasses as many source and target entities as possible, $(ii)$ entities and relations within $\mathcal{G}_{\text{sub}}$ exhibit a strong relevance to the question, and $(iii)$ $\mathcal{G}_{\text{sub}}$ is concise with little redundant information such that it can be fed into LLMs with limited lengths.
However, it is challenging to find such a $\mathcal{G}_{\text{sub}}$ since extracting a subgraph is NP-hard. To effectively and efficiently find a satisfactory $\mathcal{G}_{\text{sub}}$, we develop a tailored knowledge extraction method, named $\mathcal{P}_{\text{RL}}$, that employs deep RL to sample reasoning paths in a trial-and-error fashion. Specifically, we assume $\mathcal{G}_{\text{sub}}$ is constructed based on a set of reasoning paths $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$, where each path $\mathcal{P}_{i}=\{(e_{i},r_{1},t_{1}),(t_{1},r_{2},t_{2}),...,(t_{|\mathcal{P}_{i}|-1},r_{|\mathcal{P}_{i}|},t_{|\mathcal{P}_{i}|})\}$ is a path in $\mathcal{G}$ starting from the $i$-$th$ source entity in $\mathcal{Q}_{s}$, and $|\mathcal{P}_{i}|$ is the path length. $\mathcal{G}_{\text{sub}}$ encompasses all the entities and relations appeared in $\mathcal{P}$. We model the sampling of each reasoning path as a Markov Decision Process (MDP) with state, action, transition, and reward, defined as follows.
  • State: A state indicates the current location in KG, i.e., one of the entities in KG. Specifically, it represents the spatial change from entity $h$ to $t$. Inspired by the prior study Xiong et al. (2017), we define the state vector $\boldsymbol{s}$ as:
@@
    where $\boldsymbol{e}_{t}$ and $\boldsymbol{e}_{target}$ are the embedding vectors of the current entity and the target entity. To get the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study Feng et al. (2020). Specifically, we transform knowledge from the KG into sentences and feed them into pre-trained LM to get node embeddings.
@@
[…]
@@
Reaching a target entity is not our sole focus. To avoid overlong and rigmarole paths, we also design two auxiliary rewards to promote context relatedness and path conciseness.

### v5

# KnowGPT: Knowledge Graph based PrompTing for Large Language Models

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in many real-world applications. Nonetheless, LLMs are often criticized for their tendency to produce hallucinations, wherein the models fabricate incorrect statements on tasks beyond their knowledge and perception. To alleviate this issue, researchers have explored leveraging the factual knowledge in knowledge graphs (KGs) to ground the LLM’s responses in established facts and principles. However, most state-of-the-art LLMs are closed-source, making it challenging to develop a prompting framework that can efficiently and effectively integrate KGs into LLMs with hard prompts only. Generally, it usually suffers from three critical issues, including huge search space, high API costs, and laborious prompt engineering, that impede the widespread application in practice. To this end, we introduce a novel Knowledge Graph-based PrompTing framework, namely KnowGPT, to enhance LLMs with domain knowledge. KnowGPT contains a knowledge extraction module to extract the most informative knowledge from KGs, and a context-aware prompt construction module to automatically convert extracted knowledge into effective prompts. Experiments on three benchmark datasets demonstrate that KnowGPT significantly outperforms all competitors. Notably, KnowGPT achieves a $92.6\%$ accuracy on OpenbookQA leaderboard, close to human-level performance.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 [43] and Claude 3 [3], have surprised the world with superior performance in a wide range of real-world applications [27; 71]. Despite their impressive performance, LLMs are frequently criticized for their limited ability to handle factual information accurately and their tendency to generate hallucinations, especially when faced with questions requiring domain-specific or professional knowledge not covered in their training corpus [2; 48]. For example, when queried about nutrient composition, an LLM might erroneously associate it with “energy”, as depicted in Figure 1. This error stems from the model’s insufficient biological knowledge. Therefore, integrating domain knowledge into LLMs is crucial for reducing hallucinations and unlocking their full potential in diverse industry applications [17].
Recently, retrieval-augmented generation has been explored, which can enhance LLMs with external knowledge from text corpora or online sources [30; 76; 42]. It combines LLMs with external knowledge retrieval systems to help reduce hallucinations. However, these models face challenges in real-world applications due to the varying quality of available data. Domain knowledge is often scattered across different sources, such as textbooks, research papers, technical manuals, and industry reports [31]. These textual documents may have varying levels of quality, accuracy, and completeness, leading to potential inconsistencies or errors in the retrieved knowledge [78].
@@
Figure 1: A real-world question from OpenbookQA. GPT-3.5 could effectively correct the answer given the scientific reasoning background from ConceptNet (blue: question concepts, red: answers, grey: entities not present in questions).
A promising avenue for addressing the above issue entails the integration of Knowledge Graphs (KGs) into LLMs. KGs provide a structured representation of domain knowledge, as they are constructed based on rigid ontologies that clearly define the jargon, acronyms, specialized terminologies and their relationships in specific domains. The enormous factual knowledge stored in KGs holds the potential to ground the model’s responses in established facts and principles [20; 66; 45]. For instance, in Figure 1, an LLM can correct itself by leveraging the related background knowledge in ConceptNet [51]. Earlier studies adopted a heuristic way to inject knowledge from KGs into the LLMs during pre-training or fine-tuning. For example, ERNIE [52] incorporates entity embeddings and aligns them with word embeddings in the pre-training phase, encouraging the model to better understand and reason over entities. Similarly, KnowBERT [46] integrates entity linkers with BERT to infuse knowledge about entities during fine-tuning for knowledge-intensive applications. Another line of work focuses on retrieving relevant knowledge from KGs at inference time to augment the language model’s context. Typically, K-BERT [37] uses an attention mechanism to select relevant triples from KGs based on the query context, which are then appended to the input sequence. Similarly, KEPLER [62] learns joint embeddings of text and KG entities to enhance the model’s predictions. Subsequent works have further integrated graph neural networks alongside LLMs for joint reasoning [69; 74; 10] and introduced interactions between text tokens and KG entities within the intermediate layers of LLMs [53; 56].
However, as LLMs have been keeping evolving, most SOTA LLMs remain closed-source in practice. For instance, GPT-4 [43] and Claude 3 [3] exclusively grant access through their APIs, which means we can only retrieve model responses by submitting textual inputs, with model specifics inaccessible. As such, the research focus has recently shifted towards KG prompting that enhances fixed LLMs with KG-based hard prompts [36; 45]. KG Prompting for LLMs has been a new learning paradigm in natural language processing. Specifically, CoK [60] introduces Chain-of-Knowledge prompting to decompose LLM-generated reasoning chains into evidence triples, verifying their factuality and faithfulness using an external KG. Mindmap [63] provides more transparency on LLMs’ decision-making by enabling comprehension and reasoning over structured KG inputs. RoG [39] presents a planning-retrieval-reasoning framework that synergizes LLMs and KGs for more transparent and interpretable reasoning. KGR [18] proposes an autonomous approach to retrofit LLM-generated responses by leveraging KGs to extract, verify, and refine factual information throughout the reasoning process, effectively mitigating hallucination for knowledge-intensive applications.
Despite the promising performance of existing KG prompting methods, three critical issues hinder their widespread application in practice. Huge search space. Real-world KGs often consist of millions of triples, resulting in a vast search space when retrieving relevant knowledge for prompting. High API cost. Closed-source LLMs, like GPT-4 and Claude 3, are accessible through proprietary APIs, which can incur significant costs when performing KG prompting at scale. Thus, careful selection of the most informative knowledge from KGs is essential to minimize costs. Laborious prompt design. LLMs are highly sensitive to prompts, with even minor variations in prompts conveying the same semantic meaning potentially yielding drastically different responses. However, existing methods rely on manually designed or rule-based prompts to present factual knowledge from KGs. These hard prompts are inherently inflexible and rigid, lacking the adaptability to accommodate variations in question semantics and KG structures.
To this end, we propose a novel Knowledge Graph-based PrompTing framework, namely KnowGPT, which leverages the factual knowledge in KGs to ground the model’s responses in established facts and principles. In this paper, we aim to answer two key research questions. Given a query and a large-scale KG, how could we effectively and efficiently retrieve factual knowledge from KG that is relevant to the query? Given the raw knowledge extracted from KGs, how could we convert the extracted knowledge into an effective prompt that is easily understandable for LLM?
@@
We shed light on the the above questions with a novel prompt learning method that is rather effective, generalizable, and cost-efficient. Specifically, to address question , we leverage deep reinforcement learning (RL) to extract the most informative knowledge from KGs. To encourage the agent to discover more informative knowledge chains, we devise a tailored reward scheme that promotes the reachability, context-relatedness, and conciseness of the extracted paths. Then, a policy network is trained to maximize the reward using training questions and applied to unseen questions. To tackle question , we introduce a prompt construction strategy based on Multi-Armed Bandit (MAB). Given several knowledge extraction strategies and prompt templates, an MAB is learned to select the most effective combination for each question by balancing exploration and exploitation. The learned MAB is then applied to new questions to select knowledge extraction strategies and prompt templates automatically. Our main contributions are summarized as follows:
  • Formally define the problem of KG-based prompting, which leverages the structured knowledge in KGs to ground the LLM’s responses in established facts and principles.
  • Propose KnowGPT, a novel prompting framework that leverages deep reinforcement learning (RL) and Multi-Armed Bandit (MAB) to generate effective prompts for domain-specific queries.
  • Implement KnowGPT upon GPT-3.5. Experiments on three QA datasets shows KnowGPT outperforms SOTA baseline models by a large margin. Notably, KnowGPT achieves an average improvement of 23.7% over GPT-3.5 and an average improvement of 2.9% over GPT-4. Additionally, it attains a 92.6% accuracy on the OpenbookQA leaderboard, which is comparable to human performance.
@@ 2 Problem Statement
We formally define the problem of knowledge graph based prompting for LLMs in question answering. We represent each question as a question context $\mathcal{Q}=\{\mathcal{Q}_{s},\mathcal{Q}_{t}\}$, where $\mathcal{Q}_{s}=\{e_{1},e_{2},...,e_{m}\}$ is a set of $m$ source entities, and $\mathcal{Q}_{t}=\{e_{1},e_{2},...,e_{n}\}$ is a set of $n$ target entities. Following prior work [13; 68], $\mathcal{Q}_{s}$ is extracted by concept recognition, and we assume it is given in our problem. Similarly, each target entity in $\mathcal{Q}_{t}$ is extracted from a corresponding candidate answer. We denote an LLM as $f$, a real-world KG as $\mathcal{G}$, which consists of triples (head entity, relation, tail entity), denoted as $(h,r,t)$. In our setting, we only have access to the APIs of $f$. However, we can employ open-source lightweight language models (not $f$), like Bert-Base [24], to initialize question embeddings. Using the above notations, we describe our problem below.
Given a question $\mathcal{Q}$, an LLM $f$, and a domain KG $\mathcal{G}$, we aim to learn a prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$, which generates a prompt x that incorporates the context of $\mathcal{Q}$ and the factual knowledge in $\mathcal{G}$, such that the prediction of the LLM $f(\textbf{x})$ can output the correct answers for $\mathcal{Q}$.
Figure 2: The overall architecture of our proposed knowledge graph prompting framework, i.e., KnowGPT. Given the question context with multiple choices, we first retrieve a question-specific subgraph from the real-world KG. Knowledge Extraction is first dedicated to searching for the most informative and concise reasoning background subject to the context. Then the Prompt Construction module is optimized to prioritize the combination of knowledge and formats subject to the given question.
@@ 3 KnowGPT Framework
Learning the prompting function $f_{\text{prompt}}(\mathcal{Q},\mathcal{G})$ involves two challenges, i.e., what knowledge should be used in $\mathcal{G}$, and how to construct the prompt. To address these challenges, we present KnowGPT, which extracts raw knowledge with deep RL and then constructs the prompt with MAB. An overview of our framework is shown in Figure 2.
@@ 3.1 Knowledge Extraction with Deep Reinforcement Learning
Intuitively, the relevant reasoning background lies in a question-specific subgraph $\mathcal{G}_{\text{sub}}$ that contains all the source entities $\mathcal{Q}_{s}$, target entities $\mathcal{Q}_{t}$, and their neighbors. An ideal subgraph $\mathcal{G}_{\text{sub}}$ is expected to have the following properties: $(i)$ $\mathcal{G}_{\text{sub}}$ encompasses as many source and target entities as possible, $(ii)$ the entities and relations within $\mathcal{G}_{\text{sub}}$ exhibit a strong relevance to question context, and $(iii)$ $\mathcal{G}_{\text{sub}}$ is concise with little redundant information such that it can be fed into LLMs with limited lengths.
However, it is challenging to find such a $\mathcal{G}_{\text{sub}}$ since extracting a subgraph is NP-hard. To effectively and efficiently find a satisfactory $\mathcal{G}_{\text{sub}}$, we develop a tailored knowledge extraction method, named $\mathcal{P}_{\text{RL}}$, that employs deep RL to sample reasoning chains in a trial-and-error fashion. Specifically, we assume $\mathcal{G}_{\text{sub}}$ is constructed based on a set of reasoning chains $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{m}\}$, where each knowledge chain $\mathcal{P}_{i}=\{(e_{i},r_{1},t_{1}),(t_{1},r_{2},t_{2}),...,(t_{|\mathcal{P}_{i}|-1},r_{|\mathcal{P}_{i}|},t_{|\mathcal{P}_{i}|})\}$ is a path in $\mathcal{G}$ starting from the $i$-$th$ source entity in $\mathcal{Q}_{s}$, and $|\mathcal{P}_{i}|$ is the path length. $\mathcal{G}_{\text{sub}}$ encompasses all the entities and relations appeared in $\mathcal{P}$. We model the sampling of each reasoning path as a Markov Decision Process (MDP) with state, action, transition, and reward, defined as follows.
  • State: A state indicates the current location in KG, i.e., one of the entities in KG. Specifically, it represents the spatial change from entity $h$ to $t$. Inspired by the prior study [65], we define the state vector $\boldsymbol{s}$ as:
@@
    where $\boldsymbol{e}_{t}$ and $\boldsymbol{e}_{target}$ are the embedding vectors of the current entity and the target entity. To get the initial node embeddings for entities extracted from the background KG, we adopt the approach proposed by the previous study [13]. Specifically, we transform knowledge triples from the KG into sentences and feed them into pre-trained LM to get node embeddings.
@@
    $r_{reach}=\begin{cases}+1,&\textit{if }target;\\ -1,&\textit{otherwise},\end{cases}$ (2)
@@
Reaching a target entity is not our sole focus. To avoid overlong and rigmarole reasoning chains, we also design two auxiliary rewards to promote context-relatedness and path conciseness.